the length of a rectangular garden is 184 feet less than 3 times the width. The area of the garden is 12,875 square feet. what is the perimeter of the garden?
Answer by jorel1380(3719)

First, factor the area:
w(3w-184)=12875
3w^2-184w-12875=0
Factoring, we get:
(3w+125)(w-103)=0
w=-125/3,103
so w is 103, l=3w-184=125
P=2(103+125)=456ft.!!!!!!!!!!!!!!